40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 22 19:15:22 2011, kiwamu, Update, Green Locking, rearrangement of PSL green optics 
    Reply  Fri Dec 23 01:07:41 2011, kiwamu, Update, Green Locking, rearrangement of PSL green optics part II newLayout.png
Message ID: 6145     Entry time: Thu Dec 22 19:15:22 2011     Reply to this: 6147
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: rearrangement of PSL green optics 
 As planed (#6143), rearrangement of the PSL green setup has begun.
It required to move approximately half of the green optics on the PSL table
and I finished displacing and installing the necessary optics coarsely.
So far I just have recovered the Y arm beat-note between the PSL green light.
 
 I will do a fine alignment of the X arm path on the PSL table and try obtaining the X arm beat-note tonight.
ELOG V3.1.3-